The sum of the ages of Mr. Santos and Ms. Cruz is 55. The difference is 3. Write a system of equations.

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52781)

.

The equations are as follows


    S + C = 55     (1)     (the sum of ages)

    S - C =  3     (2)     (the difference of ages, assuming that Mr. Santos is older than Ms. Cruz)

Solved, answered and explained.



Answer by Alan3354(69443)

The sum of the ages of Mr. Santos and Ms. Cruz is 55. The difference is 3. Write a system of equations.
=================
S + C = 55
S - C = 3
--------------------------
Or,
C + S = 55
C - S = 3
========================
Who is older was not specified.
Either way, one is 29, the other is 26.